OpenStudy (anonymous):

Find vector equation for the line with the following cartesian equation. b) x+3y=7

OpenStudy (anonymous):

1; 3

OpenStudy (anonymous):

r = (1,2) + t(0,1) is the answer..... but why?

OpenStudy (anonymous):

r = ( 1,2) + t(1 ; 3) ?

OpenStudy (anonymous):

i cant tell

OpenStudy (anonymous):

sorry, it's (1,2) + t(3, -1)

OpenStudy (anonymous):

yes, let point in line is ( 1; 2) , after let vertices caster : (1; 3 ) => ( 3; -1) or ( -3; 1)

OpenStudy (anonymous):

So, how do you do this in steps? W/o knowing the answer?

OpenStudy (anonymous):

x=1 , 1 + 3y = 7 => y =2 . have M(1; 2) coefficient x : 1 ; coefficient y : 3 = > vertices perpendicular ( 1; 3) => vertices parallel ( -3, 1) ==> r = ( 1;2) + t(-3;1) ok?????

OpenStudy (anonymous):

or ==> r = (1, 2) + t(3; -1)

OpenStudy (anonymous):

hmm. ok